# Rigidity results for initial data sets satisfying the dominant energy condition

### v3

@@ 3. Spin geometry for initial data sets
$\big{|}\langle(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes 1)s,s\rangle\big{|}$ $=\big{|}(q(\nu,\nu)-\mathrm{tr}(q))\,\langle(\gamma(\nu)\otimes 1)s,s\rangle\big{|}$
@@
Inserting (3.6) and (3.7) into Corollary 3.4 concludes the proof. ∎
@@
Suppose that $M$ has odd dimension $n\geq 3$ and that the map $N\colon\Sigma\to S^{n-1}$ has positive degree. Then there exists a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
@@
Moreover, $\chi s=s$ holds at each point on the boundary $\Sigma$.
Proof.
By Proposition 2.2, the operator $\mathcal{D}$ subject to the boundary conditions $\chi s=s$ has positive index. By deformation invariance of the Fredholm index, the same holds for the operator $\mathcal{D}-\frac{i}{2}\mathrm{tr}(q)$. Hence we can find a non-trivial section $s$ of $\mathcal{E}$ such that $\mathcal{D}s-\frac{i}{2}\,\mathrm{tr}(q)\,s=0$ on $M$ and $\chi s=s$ along $\Sigma$. By elliptic regularity, $s$ is smooth. Using Proposition 3.7, we obtain
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq-\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes 1)s,s\rangle$
$\qquad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
@@
  • For each $t\in S^{1}$, the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is either Lipschitz continuous with a Lipschitz constant strictly less than $1$, or else $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$, $x\mapsto(x,0)$.
@@
Fix a smooth $2$-Lipschitz function $\beta\colon[-\pi,\pi]\to[-\pi,\pi]$ such that $\beta(t)=-\pi$ for $t$ near $-\pi$, $\beta(t)=0$ for $t$ near 00, and $\beta(t)=\pi$ for $t$ near $\pi$. Moreover, let us fix a point $a\in S^{n-1}$. We consider the map
$S^{n-1}\times[-\pi,\pi]\to S^{n},\,(x,t)\mapsto\begin{cases}(\sin\beta(t)\,a,\cos\beta(t))&\text{\rm for $t\in[-\pi,0]$},\\ (\sin\beta(t)\,x,\cos\beta(t))&\text{\rm for $t\in[0,\pi]$.}\end{cases}$
This gives a map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ with the required properties. ∎
In the following, we fix a smooth map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ as in Lemma 3.9. Moreover, $r$ will denote a parameter which will be chosen sufficiently large. We consider the product $\tilde{M}=M\times S^{1}$, equipped with the metric $\tilde{g}=g+r^{2}\,dt^{2}$. We denote by $\tilde{q}$ the pull-back $(0,2)$-tensor field of $q$ under the canonical projection $\tilde{M}\to M$. The boundary of $\tilde{M}$ can be identified with $\tilde{\Sigma}=\Sigma\times S^{1}$. We define a map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ by $\tilde{N}(x,t)=\varphi(N(x),t)$ for $x\in\Sigma$ and $t\in S^{1}$. Since the map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ is of degree $1$, the degree of the map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ agrees with the degree of the map $N\colon\Sigma\to S^{n-1}$.
@@
The trace norms of $dN$ and $d\tilde{N}$ are related by
@@
By Lemma 3.9, we have $\varphi^{*}g_{S^{n}}\leq g_{S^{n-1}}+4g_{S^{1}}$. Using the definition of $\tilde{N}$, we obtain $\tilde{N}^{*}g_{S^{n}}\leq N^{*}g_{S^{n-1}}+4g_{S^{1}}$. Let us denote the singular values of $N\colon T_{x}\Sigma\to T_{N(x)}S^{n-1}$ by $\lambda_{1},\ldots,\lambda_{n-1}$. Then the eigenvalues of $N^{*}g_{S^{n-1}}$ with respect to the metric $g$ are given by $\lambda_{1}^{2},\ldots,\lambda_{n-1}^{2}$. Hence, the eigenvalues of $N^{*}g_{S^{n-1}}+4g_{S^{1}}$ with respect to the metric $\tilde{g}$ are given by $\lambda_{1}^{2},\ldots,\lambda_{n-1}^{2},4r^{-2}$. Consequently,
@@
We consider the bundles $\mathcal{S}$ and $\mathcal{E}$ over $M$ defined above. Let $\tilde{\mathcal{S}}$ denote the pull-back of $\mathcal{S}$ under the canonical projection $\tilde{M}\to M$. Similarly, let $\tilde{\mathcal{E}}$ denote the pull-back of $\mathcal{E}$ under the canonical projection $\tilde{M}\to M$. Note that $\tilde{\mathcal{S}}$ and $\tilde{\mathcal{E}}$ are bundles over $\tilde{M}$. The map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ induces a chirality operator $\tilde{\chi}$. We define $\tilde{\mu}$ and $\tilde{J}$ as in (3.1) and (3.2) on $\tilde{M}$ using $\tilde{g}$ and $\tilde{q}$. Then $\tilde{\mu}$ and $\tilde{J}$ are the pull-backs of $\mu$ and $J$, respectively.
@@
Suppose that $M$ has even dimension $n\geq 2$ and that the map $N\colon\Sigma\to S^{n-1}$ has positive degree. For each $r>0$, we can find an element $\hat{t}\in S^{1}$ and a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
@@
Moreover, we have $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ along $\Sigma$. Here, the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
@@
Since the map $N\colon\Sigma\to S^{n-1}$ has positive degree, it follows that the map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ has positive degree as well. By Proposition 3.8, we can find a section $\tilde{s}\in C^{\infty}(\tilde{M},\tilde{\mathcal{E}})$ such that
$\int_{\tilde{M}}|\tilde{s}|^{2}>0$
and
$\int_{\tilde{M}}|\nabla^{\tilde{\mathcal{E}},\tilde{q}}\tilde{s}|^{2}$ $\leq\tfrac{1}{2}\int_{\tilde{M}}(-\tilde{\mu}+|\tilde{J}|)|s|^{2}$
$\quad+\tfrac{1}{2}\int_{\tilde{\Sigma}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
Moreover, $\tilde{s}$ satisfies $\tilde{\chi}\tilde{s}=\tilde{s}$ along the boundary $\tilde{\Sigma}$. Consequently, there exists an element $\hat{t}\in S^{1}$ such that
$\int_{M\times\{\hat{t}\}}|\tilde{s}|^{2}>0$
and
@@
$\quad+\tfrac{1}{2}\int_{\tilde{\Sigma}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
It follows from Lemma 3.10 that $\|d\tilde{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}+2r^{-1}$. Restricting $\tilde{s}$ to $M\times\{\hat{t}\}$ yields a section $s\in C^{\infty}(M,\mathcal{E})$ such that
$\int_{M}|s|^{2}>0$
and
@@
Moreover, since $\tilde{\chi}\tilde{s}=\tilde{s}$ along the boundary $\tilde{\Sigma}$, we conclude that $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ at each point on $\Sigma$, where the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$. ∎
@@
Suppose that $M$ has even dimension $n\geq 2$ and that the map $N\colon\Sigma\to S^{n-1}$ has positive degree. We can find an element $\hat{t}\in S^{1}$ and a non-trivial section $s\in H^{1}(M,\mathcal{E})$ such that
Moreover, we have $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ along $\Sigma$. Here, the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
@@
Consider an arbitrary sequence $r_{l}\to\infty$. For each $l$, Proposition 3.11 implies the existence of a section $s^{(l)}\in C^{\infty}(M,\mathcal{E})$ such that
$\int_{M}|s^{(l)}|^{2}=1$
and
@@
Moreover, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{t}_{l})$ along $\Sigma$, where the map $\hat{N}_{l}:\Sigma\to S^{n}$ is defined by $\hat{N}_{l}(x)=\tilde{N}(x,\hat{t}_{l})=\varphi(N(x),\hat{t}_{l})$ for all $x\in\Sigma$. This implies
@@
After passing to a subsequence if necessary, we may assume that $\hat{t}_{l}\to\hat{t}$ for some $\hat{t}\in S^{1}$. Moreover, we may assume that the sequence $s^{(l)}$ converges weakly in $H^{1}(M,\mathcal{E})$ to some limit $s\in H^{1}(M,\mathcal{E})$. Then $s_{l}\to s$ strongly in $L^{2}(M,\mathcal{E})$. Moreover, $s_{l}|_{\Sigma}\to s|_{\Sigma}$ strongly in $L^{2}(\Sigma,\mathcal{E}|_{\Sigma})$. Thus, we conclude that
$\int_{M}|s|^{2}=1$
and
@@
Finally, on the boundary $\Sigma$, we have $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}_{l})$ for each $l$. Passing to the limit as $l\to\infty$, we conclude that $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$, where the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$. ∎
@@ 4. Proof of Theorem I
Let $M$, $\Sigma$, $g$, $q$, $N$, $H$ satisfy the assumptions in Theorem I. In particular, we assume that $\mu\geq|J|$ at each point in $M$ and $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ at each point on the boundary $\Sigma$.
Before we prove Theorem I, we prove some preliminary statements in Proposition 4.1.
@@
The following assertions hold.
@@
    At each point $p\in\Sigma$, the eigenvalues of the second fundamental form $h$ agree with the singular values of $dN_{p}:T_{p}\Sigma\to T_{N(p)}S^{n-1}$. In particular, the eigenvalues of $h$ are nonnegative.
@@
    $q^{\rm tan}=0$ along $\Sigma$.
@@
Suppose that $n\geq 3$ is an odd integer. Since by assumption $\mu\geq|J|$ on $M$ and $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ along $\Sigma$, Proposition 3.8 yields non-trivial $\nabla^{\mathcal{E},q}$-parallel section $s\in C^{\infty}(M,\mathcal{E})$ such that $\chi s=s$ holds along $\Sigma$.
@@
Since $s$ is $\nabla^{\mathcal{E},q}$-parallel,
@@
on $\Sigma$. In the next step, we differentiate (4.2) along a vector field $X\in C^{\infty}(\Sigma,T\Sigma)$. Using (4.1) we find for the left hand side
@@
$\nabla_{X}^{\mathcal{E}}(s\circ\gamma_{0}(N))$ $=(\nabla_{X}^{\mathcal{S}}s)\circ\gamma_{0}(N)+s\circ\gamma_{0}(dN(X))$
@@
For the last equality, we have used the fact that the normal component of $q(X)$ commutes with $\gamma(\nu)$ while the tangential component of $q(X)$ anticommutes with $\gamma(\nu)$. ∎
Lemma 4.3.
We have
$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}=\|dN\|_{\mathrm{tr}}\,|s|^{2}$
at each point on $\Sigma$.
Proof.
[…]
$=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}.$
This implies
$H\,|s|^{2}+i\,(\mathrm{tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}.$
By assumption, $H-|\mathrm{tr}(q)-q(\nu,\nu)|\geq\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$. Consequently,
$\|dN\|_{\mathrm{tr}}\,|s|^{2}\leq H\,|s|^{2}-|\mathrm{tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}.$
On the other hand, using Lemma 3.6, we obtain
$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}.$
Putting these facts together, the assertion follows. ∎
@@
We define a linear subspace $L\subset S_{0}$ by
$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ at each point on $M$}\}.$
@@
It follows from Lemma 4.5 that $L$ is invariant under $\gamma_{0}(N(p))$ for each point $p\in\Sigma$. Since the map $N\colon\Sigma\to S^{n-1}$ has non-zero degree, we know that the image of $N$ is all of $S^{n-1}$. Consequently, $L$ is invariant under $\gamma_{0}(v)$ for each $v\in\mathbb{R}^{n}$.
The spinor representation of the Clifford algebra $\mathbb{C}\mathrm{l}(\mathbb{R}^{n})\to\mathrm{End}(S_{0})$ is surjective. Therefore, the elements $\gamma_{0}(v)$, $v\in\mathbb{R}^{n}$, generate all of $\mathrm{End}(S_{0})$. Putting these facts together, we conclude that $L$ is invariant under $\mathrm{End}(S_{0})$. Therefore, $L=\{0\}$ or $L=S_{0}$. The latter case is impossible as this would imply that $s$ vanishes identically. To summarize, we have shown that $L=\{0\}$. Using Lemma 4.4, we conclude that $s$ is injective at each point in $M$. For dimensional reasons, $s$ is invertible at each point in $M$. ∎
@@
Let $p$ be a point on the boundary $\Sigma$. Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$ and let $e^{0}_{1},\dots,e^{0}_{n-1}$ be an orthonormal basis of $T_{N(p)}S^{n-1}$ such that $dN(e_{j})=\lambda_{j}e^{0}_{j}$, where $\lambda_{j}\geq 0$ denote the singular values of $dN$. Then $h(e_{j})=\lambda_{j}e_{j}$ and $q(e_{j})^{\text{\rm tan}}=0$ for $j=1,\ldots,n-1$.

### v4

@@ 3. Spin geometry for initial data sets
$\big|\langle(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes\textup{id})s,s\rangle\big|$ $=\big|(q(\nu,\nu)-\mathrm{tr}(q))\,\langle(\gamma(\nu)\otimes\textup{id})s,s\rangle\big|$
@@
Using (3.3) and (3.4) together with Corollary 3.4, the assertion follows. ∎
@@
Suppose that $M$ has odd dimension $n\geq 3$ and the map $N\colon\Sigma\to S^{n-1}$ has positive degree. Then there exists a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(|J|-\mu)|s|^{2}+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
Moreover, $\chi s=s$ at each point on the boundary $\Sigma$.
Proof.
By Proposition 2.4, the operator $\mathcal{D}$ subject to the boundary conditions $\chi s=s$ has positive index. By deformation invariance of the Fredholm index, the operator $\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q)$ has positive index. Hence, we can find a non-trivial section $s$ of $\mathcal{E}$ such that $\mathcal{D}s-\tfrac{i}{2}\,\mathrm{tr}(q)\,s=0$ on $M$ and $\chi s=s$ along $\Sigma$. By elliptic regularity, $s$ is smooth. Using Proposition 3.6, we obtain
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq-\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes\textup{id})s,s\rangle$
@@
  • For each $t\in S^{1}$, the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is either Lipschitz continuous with a Lipschitz constant strictly less than $1$, or else the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$.
@@
Let us fix a point $a\in S^{n-1}$ and a smooth function $\beta\colon[-\pi,\pi]\to[-\pi,\pi]$ with the following properties:
  • $\beta(t)=-\pi$ for $t\in[-\pi,-\tfrac{7\pi}{8}]$.
  • $\beta(t)=0$ for $t\in[-\tfrac{\pi}{8},\tfrac{\pi}{8}]$.
  • $\beta(t)=\pi$ for $t\in[\tfrac{7\pi}{8},\pi]$.
  • $0<\beta^{\prime}(t)\leq 2$ for $t\in(-\tfrac{7\pi}{8},-\tfrac{\pi}{8})\cup(\tfrac{\pi}{8},\tfrac{7\pi}{8})$.
We consider the map
$\varphi\colon S^{n-1}\times[-\pi,\pi]\to S^{n},\,(\xi,t)\mapsto\begin{cases}(\sin\beta(t)\,a,\cos\beta(t))&\text{\rm for $t\in[-\pi,0]$},\\ (\sin\beta(t)\,\xi,\cos\beta(t))&\text{\rm for $t\in[0,\pi]$.}\end{cases}$
It is easy to see that $\varphi$ induces a smooth map from $S^{n-1}\times S^{1}$ to $S^{n}$ with degree $\pm 1$. We next compute
$\varphi^{*}g_{S^{n}}=\begin{cases}\beta^{\prime}(t)^{2}\,dt\otimes dt&\text{\rm for $t\in[-\pi,0]$},\\ \sin^{2}\beta(t)\,g_{S^{n-1}}+\beta^{\prime}(t)^{2}\,dt\otimes dt&\text{\rm for $t\in[0,\pi]$.}\end{cases}$
Since $0\leq\beta^{\prime}(t)\leq 2$ for all $t\in[-\pi,\pi]$, it follows that $\varphi^{*}g_{S^{n}}\leq g_{S^{n-1}}+4\,dt\otimes dt$.
If $t\in[-\pi,0]$, then the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is constant. If $t\in[0,\pi]$ and $\beta(t)\neq\tfrac{\pi}{2}$, then the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is Lipschitz continuous with Lipschitz constant $\sin\beta(t)<1$. If $t\in[0,\pi]$ and $\beta(t)=\tfrac{\pi}{2}$, the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n},\,\xi\mapsto(\xi,0)$. ∎
In the following, we fix a smooth map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ as in Lemma 3.8. Moreover, $r$ will denote a parameter which will be chosen sufficiently large. We consider the product $\tilde{M}=M\times S^{1}$, equipped with the metric $\tilde{g}=g+r^{2}\,dt\otimes dt$. We denote by $\tilde{q}$ the pull-back $(0,2)$-tensor field of $q$ under the canonical projection $\tilde{M}\to M$. We denote by $\tilde{\Sigma}=\Sigma\times S^{1}$ the boundary of $\tilde{M}$. We define a map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ by $\tilde{N}(x,t)=\varphi(N(x),t)$ for $x\in\Sigma$ and $t\in S^{1}$. Since the map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ is of degree $1$, the degree of the map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ agrees with the degree of the map $N\colon\Sigma\to S^{n-1}$.
@@
The trace norm of $d\tilde{N}$ with respect to the metric $\tilde{g}$ satisfies
@@
By Lemma 3.8, we have $\varphi^{*}g_{S^{n}}\leq g_{S^{n-1}}+4g_{S^{1}}$. Using the definition of $\tilde{N}$, we obtain
$\tilde{N}^{*}g_{S^{n}}\leq N^{*}g_{S^{n-1}}+4g_{S^{1}}.$
Let us denote the singular values of $dN_{x}\colon T_{x}\Sigma\to T_{N(x)}S^{n-1}$ by $\lambda_{1},\ldots,\lambda_{n-1}$. Then the eigenvalues of $N^{*}g_{S^{n-1}}$ with respect to the metric $g$ are given by $\lambda_{1}^{2},\ldots,\lambda_{n-1}^{2}$. Hence, the eigenvalues of $N^{*}g_{S^{n-1}}+4g_{S^{1}}$ with respect to the metric $\tilde{g}$ are given by $\lambda_{1}^{2},\ldots,\lambda_{n-1}^{2},4r^{-2}$. Consequently,
@@
We consider the bundles $\mathcal{S}$ and $\mathcal{E}$ over $M$ defined above. Let $\tilde{\mathcal{S}}$ denote the pull-back of $\mathcal{S}$ under the canonical projection $\tilde{M}\to M$. Similarly, let $\tilde{\mathcal{E}}$ denote the pull-back of $\mathcal{E}$ under the canonical projection $\tilde{M}\to M$. Note that $\tilde{\mathcal{S}}$ and $\tilde{\mathcal{E}}$ are bundles over $\tilde{M}$. Since $n$ is even, the map $\gamma_{0}\colon\mathbb{R}^{n}\to\text{\rm End}(S_{0})$ can be extended to a map $\gamma_{0}\colon\mathbb{R}^{n+1}\to\text{\rm End}(S_{0})$ satisfying the Clifford relations. With this understood, the map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ induces a chirality operator $\tilde{\chi}$. We define $\tilde{\mu}$ and $\tilde{J}$ as in (3.1) and (3.2) on $\tilde{M}$ using $\tilde{g}$ and $\tilde{q}$. Then $\tilde{\mu}$ and $\tilde{J}$ are the pull-backs of $\mu$ and $J$, respectively.
@@
Suppose $M$ has even dimension $n\geq 2$ and the map $N\colon\Sigma\to S^{n-1}$ has positive degree. For each $r>0$, we can find a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
@@
[…]
@@
[…]
$\int_{\tilde{M}}|\nabla^{\tilde{\mathcal{E}},\tilde{q}}\tilde{s}|^{2}$ $\leq\tfrac{1}{2}\int_{\tilde{M}}(-\tilde{\mu}+|\tilde{J}|)|s|^{2}+\tfrac{1}{2}\int_{\tilde{\Sigma}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
Moreover, $\tilde{s}$ satisfies $\tilde{\chi}\tilde{s}=\tilde{s}$ along the boundary $\tilde{\Sigma}$. Consequently, there exists an element $\hat{t}\in S^{1}$ such that $\int_{M\times\{\hat{t}\}}|\tilde{s}|^{2}>0$ and
@@
$\quad+\tfrac{1}{2}\int_{\Sigma\times\{\hat{t}\}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
It follows from Lemma 3.9 that $\|d\tilde{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}+2r^{-1}$. Restricting $\tilde{s}$ to $M\times\{\hat{t}\}$ yields a section $s\in C^{\infty}(M,\mathcal{E})$ such that $\int_{M}|s|^{2}>0$ and
@@
[…]
@@
Suppose that $M$ has even dimension $n\geq 2$ and the map $N\colon\Sigma\to S^{n-1}$ has positive degree. We can find a non-trivial section $s\in H^{1}(M,\mathcal{E})$ such that
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
[…]
@@
Consider an arbitrary sequence $r_{l}\to\infty$. For each $l$, Proposition 3.10 implies the existence of a section $s^{(l)}\in C^{\infty}(M,\mathcal{E})$ such that $\int_{M}|s^{(l)}|^{2}=1$ and
@@
Moreover, for each $l$, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}_{l})$ along $\Sigma$, where the map $\hat{N}_{l}\colon\Sigma\to S^{n}$ is defined by $\hat{N}_{l}(x)=\tilde{N}(x,\hat{t}_{l})=\varphi(N(x),\hat{t}_{l})$ for all $x\in\Sigma$. This implies
@@
After passing to a subsequence if necessary, we may assume that $\hat{t}_{l}\to\hat{t}$ for some $\hat{t}\in S^{1}$. Moreover, we may assume that the sequence $s^{(l)}$ converges weakly in $H^{1}(M,\mathcal{E})$ to some limit $s\in H^{1}(M,\mathcal{E})$. Then $s_{l}\to s$ strongly in $L^{2}(M,\mathcal{E})$. Moreover, $s_{l}|_{\Sigma}\to s|_{\Sigma}$ strongly in $L^{2}(\Sigma,\mathcal{E}|_{\Sigma})$. Thus, we conclude that $\int_{M}|s|^{2}=1$ and
@@
[…]
@@ 4. Proof of Theorem I
Throughout this section, we consider an initial data set $(M,g,q)$ together with a map $N\colon\Sigma\to S^{n-1}$ satisfying the assumptions in Theorem I. In particular, we assume that $\mu\geq|J|$ at each point in $M$ and $H\geq\|dN\|_{\mathrm{tr}}+|\mathrm{tr}(q)-q(\nu,\nu)|$ at each point on the boundary $\Sigma$.
@@
The following statements hold.
@@
    At each point $p\in\Sigma$, the eigenvalues of the second fundamental form $h$ agree with the singular values of $dN_{p}\colon T_{p}\Sigma\to T_{N(p)}S^{n-1}$. In particular, the eigenvalues of $h$ are nonnegative.
@@
    We have $q^{\rm tan}=0$ at each point on $\Sigma$.
@@
Suppose that $n\geq 3$ is an odd integer. By assumption, we have $\mu\geq|J|$ on $M$ and $H\geq\|dN\|_{\mathrm{tr}}+|\mathrm{tr}(q)-q(\nu,\nu)|$ along $\Sigma$. Therefore, Proposition 3.7 yields non-trivial $\nabla^{\mathcal{E},q}$-parallel section $s\in C^{\infty}(M,\mathcal{E})$ such that $\chi s=s$ along $\Sigma$.
@@
Since $s$ is $\nabla^{\mathcal{E},q}$-parallel, we obtain
@@
on $\Sigma$. In the next step, we fix a tangential vector field $X$ along $\Sigma$ and differentiate (4.2) along $X$. Using (4.1) we find for the left hand side
@@
$\nabla_{X}^{\mathcal{E}}(s\circ\gamma_{0}(N))$ $=(\nabla_{X}^{\mathcal{E}}s)\circ\gamma_{0}(N)+s\circ\gamma_{0}(dN(X))$
@@
To justify the last equality, we decompose $q(X)$ into its tangential and normal components and observe that $\gamma(q(X)^{\text{\rm tan}})$ anticommutes with $\gamma(\nu)$. ∎
@@
We define a linear subspace $L\subset S_{0}$ by $L:=\{\sigma\in S_{0}\colon\text{\rm$s\sigma=0$ at each point in $M$}\}$.
@@
It follows from Lemma 4.4 that $L$ is invariant under $\gamma_{0}(N(p))$ for each point $p\in\Sigma$. Since the map $N\colon\Sigma\to S^{n-1}$ has non-zero degree, we know that the image of $N$ is all of $S^{n-1}$. Consequently, $L$ is invariant under $\gamma_{0}(\xi)$ for each $\xi\in\mathbb{R}^{n}$.
The spinor representation of the Clifford algebra $\mathbb{C}\mathrm{l}(\mathbb{R}^{n})\to\mathrm{End}(S_{0})$ is surjective. Therefore, the elements $\gamma_{0}(\xi)$, $\xi\in\mathbb{R}^{n}$, generate all of $\mathrm{End}(S_{0})$. Putting these facts together, we conclude that $L$ is invariant under $\mathrm{End}(S_{0})$. Therefore, $L=\{0\}$ or $L=S_{0}$. The latter case is impossible as this would imply that $s$ vanishes identically. To summarize, we have shown that $L=\{0\}$. Using Lemma 4.3, we conclude that $s$ is injective at each point in $M$. For dimensional reasons, $s$ is invertible at each point in $M$. ∎
@@
Let $p$ be a point on the boundary $\Sigma$. Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$ and $e^{0}_{1},\dots,e^{0}_{n-1}$ an orthonormal basis of $T_{N(p)}S^{n-1}$ such that $dN(e_{j})=\lambda_{j}e^{0}_{j}$, where $\lambda_{j}\geq 0$ denote the singular values of $dN$. Then $h(e_{j})=\lambda_{j}e_{j}$ and $q(e_{j})^{\text{\rm tan}}=0$ for $j=1,\ldots,n-1$.